how do i subtract or add this correctly?
(4/a-5)-(1/5-a)
My answer was 4/a+a-26/5
What is the correct answer?
Thank you so much

Answer by stanbon(75887)

You have to see that -(5-a)=a-5
If you do you see the denominators
differ only by their sign.
So, rewrite at problem as follows:
4/(a-5)-1/-(a-5)
=4/(a-5)+1/(a-5)
= 5/(a-5)
Cheers,
Stan H.